1. A person has $ 100 to spend on two goods X and Y whose respective prices are $3 and $5.

A. Draw the budget line.
B. What happens to the original budget line if the budget falls by 25%?

C. What happens to the original budget line if the price of X doubles?
D. What happens to the original budget line if the price of Y falls to $4?
2. A rational consumer spends all of her income on two goods: Apple and Banana. Suppose
the last dollar spent on Apple increased her total utility from 60 utils to 68 utils and the last
dollar spent on Banana increased her total utility from 25 utils to 29 utils. If the price of a
unit of Apple is 2 Bir, what is the price of a unit of Banana at equilibrium?
1
Expert's answer
2020-02-11T09:24:38-0500

Solution:

A.

"p_x=3, G_x=\frac {100}{3}=33\frac{1}{3}"

"p_y=5, G_y=\frac{100}{5}=20"


B.


"100-0.25\times 100=75"


"p_x=3, G_x=\frac {75}{3}=25"

"p_y=5, G_y=\frac{75}{5}=15"


C.


"p_x=6, G_x=\frac {100}{6}=16\frac{2}{3}"



D.

"p_y=5, G_y=\frac{100}{4}=25"




2.


"MU_x=68-60=8, p_x=2"


"MU_y=29-25=4, p_y-?"


"\frac {MU_x}{p_x}=\frac{MU_y}{p_y}"


"\frac{8}{2}=\frac {4}{p_y}"

"p_y=1"